LBNL TPC (Software) Meeting, 10-FEB-2000, 1 pm (PST), Room 70-171

   LBNL: I.Sakrejda (chair), R.Bossingham (minutes), D.Hardtke,
         E.Hjort, H.Matis, G.Odyniec, D.Olson, H.G.Ritter,
         R.Snellings, F.Wang, H.Wieman, N.Xu
   BNL:  B.Stringfellow (calling on the 510-486-7333 conf. line)

Main topics:
   1) Offline ASIC study (N.Xu)
   2) TPC operations during the upcoming test run at RHIC

1) Nu Xu reported progress on examining the effect of the ASIC cuts;
   future work was discussed.

2) Blair Stringfellow raised several questions about TPC operations
   during the upcoming run; see mail archives:

Geometry events

The corruption of several (inner subsector) rows of Sector 9
during the Dec. '99 run is not yet understood.  Craig Consiglio
did not find any misrun FEE cables, but found an RDO-extension
cable not full plugged in, that might have led to missing bits.
Tonko took 2 geometry events on Feb.9, from which he and Blair
found that FEE 146 was read with ID=24.  Eric and/or Roy will
also look at these events.

It was decided that, for the time being, a geometry event will
be taken every other day (and after any physical changes to the
FEE).  Eric will resurrect the semi-automated geometry event
checking in the Pad Monitor, allowing immediate turnaround.

Anode HV settings

Tentatively, we will continue to use the settings used in Dec. '99.
However, Sergei Panitkin has promised noise data broken down by
subsector, and we may reconsider when this is available.  (No major
changes are foreseen.)

ASIC settings

We have asked for 1000, full-field cosmics with ASIC=1,3,2,0
setting (for a discussion of the parameters, see:

If the data size is acceptable, these settings will be retained
for the remainder of the upcoming run.

TCD settings

Since a TCD board had to be appropriated for MWC threshold setting,
and because the variable-freq. clock showed stability problems in
Dec. '99, a fixed, 9.38-MHz ADC sampling clock will be used for
the test run, standing in for the RHIC strobe.

One laser will be phase-locked to this clock.  There is only one
board, but (if power supply problems on the second laser can be
overcome) the second laser will be run as well; their relative
jitter will be measured by Alexei.

ADC Sampling Range

Tonko has inquired about eliminating time bins <10 and >400;
it was agreed that, for the moment, we would only eliminate
those >400 (which are well beyond the central membrane).

Bad FEE Criteria?

Craig Consiglio currently plans to replace 30 FEE's on the East End
of the TPC within the next day, or so.  The failure modes are the
failed cards were not known at the meeting; Chinh Vu will be queried.
Current rejection criteria by Craig C. and Dan C. were not known
by those in attendance.

Ground-Plane Pulser

Iwona has extensive information from Frabrice Retiere regarding the
analysis of the pulser data.  Frank Laue continues to be interested
in analyzing the t0 data from the pulser, and might possibly do the
gain work, as well.

We requested a pulser run ***with the MWC cards powered*** when
it is possible; field-off is OK.

Pedestal Runs

Studies during summer, '99 by Iwona and Oleg found the result
of the first pedestal run after powering to be consistently low
by 0.6 ADC counts, but to be stable thereafter.  This is not
understood, but was repeatable.  Therefore, two pedestal runs
should be taken.  It was agreed that taking pedestals every
10,000 events, or so, was a reasonable measure until stability
is better known.

Schedule Information (from Blair)

Feb. 14-21:  MWC work will occur next week, and this will involve
             trigger and the TPC, but not DAQ.  DAQ will be available
             only for short runs (<1000 events).

Feb. 22-23:  Poletips go in, shielding wall goes up.

Feb. 24:     Magnet first available around then.

Mar. 3:      End of run

Data taking for Feb. 14-21 period (sans magnetic field)

 - Iwona & Eric requested 20K cosmics w/o field, using RICH trigger;
   these are for studying electric-field distortions.

 - Blair hopes for field-off laser run.

 - Roy requested pulser run w/MWC's powered.

Wo/Man Power

 - Iwona will be at BNL for 2 weeks.
 - Lidia can now run cosmic-ray data production jobs.
 - Eric will be there for the last week of the run.
 - Jim Thomas is back at BNL now.
 - Howard W. returns ~Feb. 17.